Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2017-09-27

Observers: Kris Rybicki, Simon Hodgkin
TAT: Regis Cartier, Janet Chen
Support team: Morgan Fraser, Emma Callis

Calibrations: biases, dome arcs+flats - apparently some problem with flats

Conditions: Thick cloud, closed at sunset.
00:45 Opened! (still somewhat cloudy)
01:17 Seeing 0.9 arcsec
~01:20 Refocussed
~01:50 Refocussed (Guide star looked like a comet)`
~03:35 Closed for Cloud
05:00 Reopening after cloud - but an issue with an interlock (dome/instrument rotation?) has forced the TO to go up to the telescope to clear it.
06:05 got on telluric STD. It's definitely clearer - though not clear - but target is hard to identify.
06:35 Swapped back to EFOSC
Then CLOSED for good, almost immediately


UT       Target            Type            Comment
==============================================================================================     
12:55    EG131            STD             Still some cloud
01:34    AT2017gwx        Classify        Looks positively murky. Spectrum blue, Z=0 Balmer series in absorption, Halpha in emission.
01:55    AT2017gwt        Classify        SN Ia (norm, close to max, Z~0.06)
02:15    AT2017gwy        Classify        SN Ia (also close to max)
02:30    ATLAS17lau       Classify        Aimed for the galaxy nucleus in hope, but nothing obvious there: just a faint galaxy spectrum
02:50    Feige 110
03:22    AT2017gww        Classify        Observed in thickening cloud with GR13. Looks like a young SN type II
06:05    HIP024337_A      Telluric
06:15    N113-YSO04_A     Acquisition     Target not obvious from finding chart
06:53    Feige 110                        Aborted

Comments